# Tri-hypercharge versus tri-darkcharge

### v1

@@ IV.3 Number analysis of fermion spectrum
$\sin\theta^{(q)}_{12}$ $0.22501\pm 0.00068$ $0.225011$
$\sin\theta^{(q)}_{23}$ $0.04183_{-0.00069}^{+0.00079}$ $0.041829$
$\sin\theta^{(q)}_{13}$ $0.003732^{+0.000090}_{-0.000085}$ $0.003732$
@@
[…]
@@
$\delta^{(\ell)}_{\text{CP}}$ $[^{\circ}]$ $212_{-41}^{+26}$ 124–364 211.995
@@ V.1 $\rho$ parameter
Using the global fit value, $\rho=1.00031\pm 0.00019$ [4], we derive a low bound on the new physics scale $v_{23}$, shown as the red curve in Fig. 5. In this analysis, we have taken $v=246$ GeV, $s_{W}^{2}=0.231$, $g=0.652$, and imposed $\Delta\rho=0.0005$. The shaded region corresponds to the excluded parameter space. The constraint is most stringent for $g_{3}\gtrsim 0.8$. For reference, the dashed blue line represents a natural gauge unification scenario in which the three gauge couplings are of similar magnitude, i.e., $g_{1}=g_{2}=g_{3}=\sqrt{3}g_{Y}$. In this case, we obtain the lower bounds $v_{23}\gtrsim 6.88$ TeV and $m_{Z_{23}}\gtrsim 2.78$ TeV.
@@ V.2 $\varepsilon_{1,2}$ mixing parameters
as explicitly shown in Table 6 in Appendix B. According to the electroweak precision data [4], the new physics effects remain consistent with observations if $|\varepsilon_{1,2}|\lesssim 10^{-3}$ [49, 50]. Hence, we take the bound $|\varepsilon_{1}|=10^{-3}$ into account and then obtain the allowed parameter space as determined by the brown curve in Fig. 5. In the gauge unification scenario $g_{1}=g_{2}=g_{3}=\sqrt{3}g_{Y}$, this translates to the bounds $v_{23}\gtrsim 5.95$ TeV and $m_{Z_{23}}\gtrsim 2.39$ TeV.
@@ VI.1 LEPII
It is checked that the strongest constraint on $m_{Z_{23}}$, and thus $v_{23}$, comes from the $VV$ model with $\Lambda_{\mu\mu}^{+}=18.9$ TeV, which is displayed by orange curve in Fig. 5. In the region $g_{3}\gtrsim 0.4$, this bound is much lower than that from the LHC discussed below. However, for $g_{Y}\lesssim g_{3}\lesssim 0.4$, the LEPII constraint is significant, implying $m_{Z_{23}}\gtrsim 6$ TeV.
@@ VI.2 LHC
The bound of parameter space induced by the process $pp\to Z^{\prime}\to l\bar{l}$ with the latest ATLAS [52] (CMS [53]) constraint taking width per resonance mass to be 3% (0.6%) is shown in Fig. 5 by a black (gray) curve. This bound is strongest for $0.4\lesssim g_{3}\lesssim 0.8$ and implies that in gauge unification scenario $g_{1}=g_{2}=g_{3}=\sqrt{3}g_{Y}$ then $v_{23}\gtrsim 12$ TeV and $m_{Z_{23}}\gtrsim 4.9$ TeV.
@@ VII Flavor constraints
Due to the non-universal charge assignments of fermion generations under the gauge group $U(1)_{Y_{1}}\otimes U(1)_{Y_{2}}\otimes U(1)_{Y_{3}}$, the model naturally contains tree-level flavor-changing neutral currents (FCNCs). These are mediated by the heavy gauge bosons $Z_{12}$ and $Z_{23}$, as well as the SM $Z$ boson, and contribute to $\Delta F=2$ processes, such as meson mass differences $\Delta m_{P}$ ($P=K,B_{s},B_{d},D$). It is important to note that the parameters $y^{d}_{i}$, which account for the observed fermion mass hierarchies and mixing, are taken to be real. Consequently, the model does not introduce new sources of CP violation in neutral meson mixing or rare decay processes. Additionally, the FCNC interactions also affect several $b\to sll$ observables and lepton flavor-violating decays. In this work, we provide a more detailed analysis of the FCNC phenomenology than in previous studies, such as Refs. [41, 42].
@@
where the flavor-violating couplings $[\Gamma^{Z_{12},Z_{23},Z}_{L(R)}]_{ij}$ induced by gauge bosons $Z_{12},Z_{23},Z$ are generally have the below forms
@@
in which $Y^{L,R}_{i}$ are hypercharges for $q_{iL,R}$, i.e $Y^{L}=\left\{1/6,1/6\right\}$ for $u_{iL},d_{iL}$, $Y^{R}=\left\{2/3,-1/3\right\}$ for $u_{iR},d_{iR}$. The unitary matrices $V_{q_{L,R}}$ connect the weak and mass eigenstates, which can be numerically obtained by benchmark points that successfully reproduce the fermion spectrum, as shown in Table 4. It should be noted that with $Y^{1}_{L}=Y^{2}_{L}=Y^{3}_{L}$, the flavor-violating couplings of SM $Z$ boson $[\Gamma^{Z}_{L(R)}]_{ij}$ vanish, due to the unitary condition of $V_{qL(R)}$, i.e $[\Gamma^{Z}_{L(R)}]_{ij}\sim\sum_{k=1}^{3}[V^{*}_{qL(R)}]_{ki}[V_{qL(R)}]_{kj}=0$ with $i\neq j$. In addition, the terms depend on $1/m^{2}_{Z_{12}}$ in the above WCs can be skipped due to a very high scale $m_{Z_{12}}\sim v_{12}\sim\mathcal{O}(10^{3})$ TeV. This leads the WCs in $\Delta F=2$ processes depend mostly on $m_{Z_{23}}$ and couplings $g_{12},g_{3}$.
@@
$[\tilde{g}^{Z_{23}}_{V,A}]_{ij}=\frac{1}{2}\sum_{k=1}^{3}\left[(V^{*}_{eL})_{ki}\left(g^{Z_{23}}_{V}(e_{k})+g^{Z_{23}}_{A}(e_{k})\right)(V_{eL})_{kj}\pm(V^{*}_{eR})_{ki}\left(g^{Z_{23}}_{V}(e_{k})-g^{Z_{23}}_{A}(e_{k})\right)(V_{eR})_{kj}\right]$ (117)
are vector and axial-vector couplings induced by new gauge boson $Z_{23}$ in mass eigenstates, while $g^{Z_{23}}_{V,A}(e_{k})$ are vector and axial-vector couplings in flavor states and given explicitly in Tables 6 and 7.
@@
×$\times\left(|[C^{\text{NP}}_{9}]_{ij}-[C^{{}^{\prime}\text{NP}}_{9}]_{ij}|^{2}+|[C_{10}^{\text{NP}}]_{ij}-[C^{{}^{\prime}\text{NP}}_{10}]_{ij}|^{2}\right).$
Among above processes with different product lepton flavors $(e\mu,e\tau,\mu\tau)$, we concentrate on the ones with decaying lepton flavors $e^{+}\mu^{-}$ since they have strongest constraint, namely BR$(B_{s}\to e^{\pm}\mu^{\mp})_{\text{exp}}<5.4\times 10^{-9}$ [4].
@@
for $i\neq j$, $i=\mu,\tau$ and $j=e,\mu$. The factor $\eta_{\text{RGE}}$ presents for RGE running from high scale $\mu\sim m_{Z_{23}}\sim\mathcal{O}(10)$ TeV to low scale $\mu\sim 1$ GeV, which numerically is $\eta_{\text{RGE}}\sim 0.9$ [42]. In addition, the couplings $[\tilde{g}^{Z_{23}}_{L(R)}]_{ij}$ are given in Eq. (117). Besides three-body leptonic decays, there exist loop contributions involving $Z_{23}$ and SM charged lepton $e_{k}$ to radiative decays $e_{i}\to e_{j}\gamma$. Their branching ratios are given in the limit $m_{e_{i}}\gg m_{e_{j}}$ and $m_{e_{k}}^{2}/m^{2}_{Z_{23}}\ll 1$ as follow
@@
with coefficients $[C^{Z_{23}}_{L(R)}]_{ij}$ read
@@
For numerical study, we focus on the observables having the strongest experimental constraints, BR$(\mu\to 3e)_{\text{exp}}<1.0\times 10^{-12}$ and BR$(\mu\to e\gamma)_{\text{exp}}<4.2\times 10^{-13}$ [4].
Figure 6: Black, brown, orange, pink, red, and purple curves denote the exclusion bounds derived from $\epsilon_{B_{s}}$, $\epsilon_{B_{d}}$, $\epsilon_{B_{s}\to\mu^{+}\mu^{-}}$, BR$(B_{s}\to e^{\pm}\mu^{\mp})$, BR$(\mu\to 3e)$, and BR$(\mu\to e\gamma)$, respectively. The shaded regions indicate the excluded parameter space. The dashed blue line corresponds to the gauge unification scenario $g_{1}=g_{2}=g_{3}=\sqrt{3}g_{Y}$.
In Fig. 6, we show the correlation between the coupling $g_{3}$ of gauge group $U(1)_{Y_{3}}$ and the VEV $v_{23}$ satisfying constraints of $\epsilon_{B_{s}}$ (black), $\epsilon_{B_{d}}$ (brown), $\epsilon_{B_{s}\to\mu^{+}\mu^{-}}$ (orange), BR$(B_{s}\to e^{\pm}\mu^{\mp})$ (pink), BR$(\mu\to 3e)$ (red), and BR$(\mu\to e\gamma)$ (purple). The parameter space, as indicated by the shaded regions, is excluded. We see that the process $\mu\to 3e$ ($\mu\to e\gamma$) gives the strongest constraint if $g_{3}\lesssim 1$ ($g_{3}\gtrsim 1$). At $g_{3}\simeq 1$ then $v_{23}$ and $m_{Z_{23}}$ are minimal, i.e., $v_{23}\simeq 11.6$ TeV and $m_{Z_{23}}\simeq 6.3$ TeV. For the gauge unification case $g_{1}=g_{2}=g_{3}=\sqrt{3}g_{Y}$, we have $v_{23}\gtrsim 18.6$ TeV and $m_{Z_{23}}\gtrsim 7.3$ TeV. These lower bounds are larger than the bounds given in sections V and VI.
@@ VIII.1 Dark matter relic abundance
×$\times\left(1-\frac{m_{\mathsf{H}}^{2}}{M_{1}^{2}}\right)^{\frac{1}{2}}\Theta(M_{1}-m_{\mathsf{H}})$
@@ VIII.2 Dark matter scattering off nuclei
with the loop function $\mathcal{G}_{2}(x)=2[x+(1-x)\ln(1-x)]/x^{2}-1$ and the couplings $g_{A}^{q}=\frac{1}{2}\left(-\frac{1}{2}\right)$ for $q=u,c,t~{}(d,s,b)$, $g_{V}^{\ell}=-\frac{g}{2c_{W}}\left(\frac{1}{2}-2s_{W}^{2}\right)$, $g_{A}^{\ell}=-\frac{g}{4c_{W}}$, $g_{V}^{\nu}=g_{A}^{\nu}=\frac{g}{4c_{W}}$. Hence, the spin-dependent (SD) scattering cross section of $N_{1R}$ in the case of the proton target is given by
@@ Appendix B Vector and axial-vector couplings
$e_{1}$ $\frac{1-2c_{2W}}{2}-\frac{3g_{1}c_{W}(c_{12}s_{23}\varepsilon_{1}+s_{12}\varepsilon_{2})}{2g}$ $-\frac{1}{2}+\frac{g_{1}c_{W}(c_{12}s_{23}\varepsilon_{1}+s_{12}\varepsilon_{2})}{2g}$
$e_{2}$ $\frac{1-2c_{2W}}{2}-\frac{3g_{2}c_{W}(s_{12}s_{23}\varepsilon_{1}-c_{12}\varepsilon_{2})}{2g}$ $-\frac{1}{2}+\frac{g_{2}c_{W}(s_{12}s_{23}\varepsilon_{1}-c_{12}\varepsilon_{2})}{2g}$
$e_{3}$ $\frac{1-2c_{2W}}{2}+\frac{3g_{3}c_{W}c_{23}\varepsilon_{1}}{2g}$ $-\frac{1}{2}-\frac{g_{3}c_{W}c_{23}\varepsilon_{1}}{2g}$
@@
$d_{1}$ $-\frac{1+2c_{2W}}{6}-\frac{g_{1}c_{W}(c_{12}s_{23}\varepsilon_{1}+s_{12}\varepsilon_{2})}{6g}$ $-\frac{1}{2}+\frac{g_{1}c_{W}(c_{12}s_{23}\varepsilon_{1}+s_{12}\varepsilon_{2})}{2g}$
$d_{2}$ $-\frac{1+2c_{2W}}{6}-\frac{g_{2}c_{W}(s_{12}s_{23}\varepsilon_{1}-c_{12}\varepsilon_{2})}{6g}$ $-\frac{1}{2}+\frac{g_{2}c_{W}(s_{12}s_{23}\varepsilon_{1}-c_{12}\varepsilon_{2})}{2g}$
$d_{3}$ $-\frac{1+2c_{2W}}{6}+\frac{g_{3}c_{W}c_{23}\varepsilon_{1}}{6g}$ $-\frac{1}{2}-\frac{g_{3}c_{W}c_{23}\varepsilon_{1}}{2g}$
@@
$\nu_{2}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $-\frac{g_{2}c_{W}c_{12}}{2g}$ $-\frac{g_{2}c_{W}c_{12}}{2g}$
$\nu_{3}$ $-\frac{g_{3}c_{W}c_{23}}{2g}$ $-\frac{g_{3}c_{W}c_{23}}{2g}$ 00 00
$e_{1}$ $\frac{3g_{12}c_{W}s_{23}}{2g}$ $-\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{3g_{1}c_{W}s_{12}}{2g}$ $-\frac{g_{1}c_{W}s_{12}}{2g}$
$e_{2}$ $\frac{3g_{12}c_{W}s_{23}}{2g}$ $-\frac{g_{12}c_{W}s_{23}}{2g}$ $-\frac{3g_{2}c_{W}c_{12}}{2g}$ $\frac{g_{2}c_{W}c_{12}}{2g}$
$e_{3}$ $-\frac{3g_{3}c_{W}c_{23}}{2g}$ $\frac{g_{3}c_{W}c_{23}}{2g}$ 00 00
$u_{1}$ $-\frac{5g_{12}c_{W}s_{23}}{6g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $-\frac{5g_{1}c_{W}s_{12}}{6g}$ $\frac{g_{1}c_{W}s_{12}}{2g}$
$u_{2}$ $-\frac{5g_{12}c_{W}s_{23}}{6g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{5g_{2}c_{W}c_{12}}{6g}$ $-\frac{g_{2}c_{W}c_{12}}{2g}$
$u_{3}$ $\frac{5g_{3}c_{W}c_{23}}{6g}$ $-\frac{g_{3}c_{W}c_{23}}{2g}$ 00 00
$d_{1}$ $\frac{g_{12}c_{W}s_{23}}{6g}$ $-\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{g_{1}c_{W}s_{12}}{6g}$ $-\frac{g_{1}c_{W}s_{12}}{2g}$
$d_{2}$ $\frac{g_{12}c_{W}s_{23}}{6g}$ $-\frac{g_{12}c_{W}s_{23}}{2g}$ $-\frac{g_{2}c_{W}c_{12}}{6g}$ $\frac{g_{2}c_{W}c_{12}}{2g}$
$d_{3}$ $-\frac{g_{3}c_{W}c_{23}}{6g}$ $\frac{g_{3}c_{W}c_{23}}{2g}$ 00 00

### v2

@@ IV.3 Number analysis of fermion spectrum
$\sin\theta^{(q)}_{12}$ $\pm 0.00068 0.2250110.2250$11
$\sin\theta^{(q)}_{23}$ $04183_{-0.00069}^{+0.00079} 0.0$418290.041829
$\sin\theta^{(q)}_{13}$ $^{+0.000090}_{-0.000085} 0.0037320.003732$
@@
$y_{11}^{u}$ $\simeq$ $-0.274808,\hskip 14.22636pty_{12}^{u}\simeq-2.8602,\hskip 14.22636pty_{13}^{u}\simeq 6.23966,\hskip 14.22636pty_{21}^{u}\simeq 1.17024,$ (63)
$y_{22}^{u}$ $\simeq$ $1.38277,\hskip 14.22636pty_{23}^{u}\simeq 0.283248,\hskip 14.22636pty_{33}^{u}\simeq 0.993968,\hskip 14.22636pt\text{arg}~(y_{13}^{u})\simeq 108.115^{\circ},$ (64)
$y_{11}^{d}$ $\simeq$ $-0.52976,\hskip 14.22636pty_{12}^{d}\simeq-1.28349,\hskip 14.22636pty_{13}^{d}\simeq 0.883176,\hskip 14.22636pty_{21}^{d}\simeq 0.598669,$ (65)
$y_{22}^{d}$ $\simeq$ $-0.52933,\hskip 14.22636pty_{23}^{d}\simeq-0.286541,\hskip 14.22636pty_{33}^{d}\simeq-0.0164254.$ (66)
@@
$\delta^{(\ell)}_{\text{CP}}$ $^{\circ}$] $212_{-41}^{+26}$ 124–364 211.995
@@ V.1 $\rho$ parameter
Using the global fit value, $\rho=1.00031\pm 0.00019$ [4], we derive a low bound on the new physics scale $v_{23}$, shown as the red curve in Fig. 5. In this analysis, we have taken $v=246$ GeV, $s_{W}^{2}=0.231$, $g=0.652$, and imposed $\Delta\rho=0.0005$. The shaded region corresponds to the excluded parameter space. The constraint is most stringent for $g_{3}\gtrsim 0.79$. For reference, the dashed blue line represents a natural gauge unification scenario in which the three gauge couplings are of similar magnitude, i.e., $g_{1}=g_{2}=g_{3}=\sqrt{3}g_{Y}$. In this case, we obtain the lower bounds $v_{23}\gtrsim 7.0$ TeV and $m_{Z_{23}}\gtrsim 2.8$ TeV.
@@ V.2 $\varepsilon_{1,2}$ mixing parameters
as explicitly shown in Table 7 in Appendix B. According to the electroweak precision data [4], the new physics effects remain consistent with observations if $|\varepsilon_{1,2}|\lesssim 10^{-3}$ [50, 51]. Hence, we take the bound $|\varepsilon_{1}|=10^{-3}$ into account and then obtain the allowed parameter space as determined by the brown curve in Fig. 5. In the gauge unification scenario $g_{1}=g_{2}=g_{3}=\sqrt{3}g_{Y}$, this translates to the bounds $v_{23}\gtrsim 6.0$ TeV and $m_{Z_{23}}\gtrsim 2.4$ TeV.
@@ VI.1 LEPII
It is checked that the strongest constraint on $m_{Z_{23}}$, and thus $v_{23}$, comes from the $VV$ model with $\Lambda_{\mu\mu}^{+}=18.9$ TeV, which is displayed by orange curve in Fig. 5. In the region $g_{3}\gtrsim 0.42$, this bound is much lower than that from the LHC discussed below. However, for $g_{Y}\lesssim g_{3}\lesssim 0.42$, the LEPII constraint is significant, implying $m_{Z_{23}}\gtrsim 6$ TeV.
@@ VI.2 LHC
The bound of parameter space induced by the process $pp\to Z^{\prime}\to l\bar{l}$ with the latest ATLAS [53] (CMS [54]) constraint taking width per resonance mass to be 3% (0.6%) is shown in Fig. 5 by a black (gray) curve. This bound is strongest for $0.42\lesssim g_{3}\lesssim 0.79$ and implies that in gauge unification scenario $g_{1}=g_{2}=g_{3}=\sqrt{3}g_{Y}$ then $v_{23}\gtrsim 12.1$ TeV and $m_{Z_{23}}\gtrsim 4.9$ TeV.
@@ VII Flavor constraints
Due to the non-universal charge assignments of fermion generations under the gauge group $U(1)_{Y_{1}}\otimes U(1)_{Y_{2}}\otimes U(1)_{Y_{3}}$, the model naturally contains tree-level flavor-changing neutral currents (FCNCs). These are mediated by the heavy gauge bosons $Z_{12}$ and $Z_{23}$, as well as the SM $Z$ boson, and contribute to $\Delta F=2$ processes, such as meson mass differences $\Delta m_{P}$ ($P=K,B_{s},B_{d},D$). It is important to note that the parameters $y^{d}_{i}$, which account for the observed fermion mass hierarchies and mixing, are taken to be real. Consequently, the model does not introduce new sources of CP violation in neutral meson mixing or rare decay processes. Additionally, the FCNC interactions also affect several $b\to sll$ observables and lepton flavor-violating decays. In this work, we provide a more detailed analysis of the FCNC phenomenology than in previous studies, such as Refs. [41, 42, 43].
@@
where the flavor-violating couplings $\Gamma^{Z_{12},Z_{23},Z}_{L(R)}]_{ij}$ induced by gauge bosons $Z_{12},Z_{23},Z$ are generally have the below forms
@@
in which $Y^{L,R}_{i}$ are hypercharges for $q_{iL,R}$, i.e $Y^{L}=\left\{1/6,1/6\right\}$ for $u_{iL},d_{iL}$, $Y^{R}=\left\{2/3,-1/3\right\}$ for $u_{iR},d_{iR}$. The unitary matrices $V_{q_{L,R}}$ connect the weak and mass eigenstates, which can be numerically obtained by benchmark points that successfully reproduce the fermion spectrum, as shown in Table 5. It should be noted that with $Y^{1}_{L}=Y^{2}_{L}=Y^{3}_{L}$, the flavor-violating couplings of SM $Z$ boson $\Gamma^{Z}_{L(R)}]_{ij}$ vanish, due to the unitary condition of $V_{qL(R)}$, i.e $\Gamma^{Z}_{L(R)}]_{ij}\sim\sum_{k=1}^{3}[V^{*}_{qL(R)}]_{ki}[V_{qL(R)}]_{kj}=0$ with $i\neq j$. In addition, the terms depend on $1/m^{2}_{Z_{12}}$ in the above WCs can be skipped due to a very high scale $m_{Z_{12}}\sim v_{12}\sim\mathcal{O}(10^{3})$ TeV. This leads the WCs in $\Delta F=2$ processes depend mostly on $m_{Z_{23}}$ and couplings $g_{12},g_{3}$.
@@
$\tilde{g}^{Z_{23}}_{V,A}]_{ij}=\frac{1}{2}\sum_{k=1}^{3}\left[(V^{*}_{eL})_{ki}\left(g^{Z_{23}}_{V}(e_{k})+g^{Z_{23}}_{A}(e_{k})\right)(V_{eL})_{kj}\pm(V^{*}_{eR})_{ki}\left(g^{Z_{23}}_{V}(e_{k})-g^{Z_{23}}_{A}(e_{k})\right)(V_{eR})_{kj}\right]$ (119)
are vector and axial-vector couplings induced by new gauge boson $Z_{23}$ in mass eigenstates, while $g^{Z_{23}}_{V,A}(e_{k})$ are vector and axial-vector couplings in flavor states and given explicitly in Tables 7 and 8.
@@
$\times\left(|[C^{\text{NP}}_{9}]_{ij}-[C^{{}^{\prime}\text{NP}}_{9}]_{ij}|^{2}+|[C_{10}^{\text{NP}}]_{ij}-[C^{{}^{\prime}\text{NP}}_{10}]_{ij}|^{2}\right).$
[…]
@@
for $i\neq j$, $i=\mu,\tau$ and $j=e,\mu$. The factor $\eta_{\text{RGE}}$ presents for RGE running from high scale $\mu\sim m_{Z_{23}}\sim\mathcal{O}(10)$ TeV to low scale $\mu\sim 1$ GeV, which numerically is $\eta_{\text{RGE}}\sim 0.9$ [42]. In addition, the couplings $\tilde{g}^{Z_{23}}_{L(R)}]_{ij}$ are given in Eq. (119). Besides three-body leptonic decays, there exist loop contributions involving $Z_{23}$ and SM charged lepton $e_{k}$ to radiative decays $e_{i}\to e_{j}\gamma$. Their branching ratios are given in the limit $m_{e_{i}}\gg m_{e_{j}}$ and $m_{e_{k}}^{2}/m^{2}_{Z_{23}}\ll 1$ as follow
@@
with coefficients $C^{Z_{23}}_{L(R)}]_{ij}$ read
@@
[…]
@@ VIII.1 Dark matter relic abundance
$\times\left(1-\frac{m_{\mathsf{H}}^{2}}{M_{1}^{2}}\right)^{\frac{1}{2}}\Theta(M_{1}-m_{\mathsf{H}})$
@@ VIII.2 Dark matter scattering off nuclei
with the loop function $\mathcal{G}_{2}(x)=2[x+(1-x)\ln(1-x)]/x^{2}-1$ and the couplings $g_{A}^{q}=\frac{1}{2}\left(-\frac{1}{2}\right)$ for q=u,c,t(d,s,b)q=u,c,t~(d,s,b), $g_{V}^{\ell}=-\frac{g}{2c_{W}}\left(\frac{1}{2}-2s_{W}^{2}\right)$, $g_{A}^{\ell}=-\frac{g}{4c_{W}}$, $g_{V}^{\nu}=g_{A}^{\nu}=\frac{g}{4c_{W}}$. Hence, the spin-dependent (SD) scattering cross section of $N_{1R}$ in the case of the proton target is given by
@@ Appendix B Vector and axial-vector couplings
$e_{1}$ $\frac{1-2c_{2W}}{2}-\frac{3g_{1}c_{W}(c_{12}s_{23}\varepsilon_{1}+s_{12}\varepsilon_{2})}{2g}$ $\frac{1}{2}+\frac{g_{1}c_{W}(c_{12}s_{23}\varepsilon_{1}+s_{12}\varepsilon_{2})}{2g}$
$e_{2}$ $\frac{1-2c_{2W}}{2}-\frac{3g_{2}c_{W}(s_{12}s_{23}\varepsilon_{1}-c_{12}\varepsilon_{2})}{2g}$ $\frac{1}{2}+\frac{g_{2}c_{W}(s_{12}s_{23}\varepsilon_{1}-c_{12}\varepsilon_{2})}{2g}$
$e_{3}$ $\frac{1-2c_{2W}}{2}+\frac{3g_{3}c_{W}c_{23}\varepsilon_{1}}{2g}$ $\frac{1}{2}-\frac{g_{3}c_{W}c_{23}\varepsilon_{1}}{2g}$
@@
$d_{1}$ $\frac{1+2c_{2W}}{6}-\frac{g_{1}c_{W}(c_{12}s_{23}\varepsilon_{1}+s_{12}\varepsilon_{2})}{6g}$ $\frac{1}{2}+\frac{g_{1}c_{W}(c_{12}s_{23}\varepsilon_{1}+s_{12}\varepsilon_{2})}{2g}$
$d_{2}$ $\frac{1+2c_{2W}}{6}-\frac{g_{2}c_{W}(s_{12}s_{23}\varepsilon_{1}-c_{12}\varepsilon_{2})}{6g}$ $\frac{1}{2}+\frac{g_{2}c_{W}(s_{12}s_{23}\varepsilon_{1}-c_{12}\varepsilon_{2})}{2g}$
$d_{3}$ $\frac{1+2c_{2W}}{6}+\frac{g_{3}c_{W}c_{23}\varepsilon_{1}}{6g}$ $\frac{1}{2}-\frac{g_{3}c_{W}c_{23}\varepsilon_{1}}{2g}$
@@
$\nu_{2}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{g_{2}c_{W}c_{12}}{2g}$ $\frac{g_{2}c_{W}c_{12}}{2g}$
$\nu_{3}$ $\frac{g_{3}c_{W}c_{23}}{2g}$ $\frac{g_{3}c_{W}c_{23}}{2g}$ 0 0
$e_{1}$ $\frac{3g_{12}c_{W}s_{23}}{2g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{3g_{1}c_{W}s_{12}}{2g}$ $\frac{g_{1}c_{W}s_{12}}{2g}$
$e_{2}$ $\frac{3g_{12}c_{W}s_{23}}{2g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{3g_{2}c_{W}c_{12}}{2g}$ $\frac{g_{2}c_{W}c_{12}}{2g}$
$e_{3}$ $\frac{3g_{3}c_{W}c_{23}}{2g}$ $\frac{g_{3}c_{W}c_{23}}{2g}$ 0 0
$u_{1}$ $\frac{5g_{12}c_{W}s_{23}}{6g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{5g_{1}c_{W}s_{12}}{6g}$ $\frac{g_{1}c_{W}s_{12}}{2g}$
$u_{2}$ $\frac{5g_{12}c_{W}s_{23}}{6g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{5g_{2}c_{W}c_{12}}{6g}$ $\frac{g_{2}c_{W}c_{12}}{2g}$
$u_{3}$ $\frac{5g_{3}c_{W}c_{23}}{6g}$ $\frac{g_{3}c_{W}c_{23}}{2g}$ 0 0
$d_{1}$ $\frac{g_{12}c_{W}s_{23}}{6g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{g_{1}c_{W}s_{12}}{6g}$ $\frac{g_{1}c_{W}s_{12}}{2g}$
$d_{2}$ $\frac{g_{12}c_{W}s_{23}}{6g}$ $\frac{g_{12}c_{W}s_{23}}{2g}$ $\frac{g_{2}c_{W}c_{12}}{6g}$ $\frac{g_{2}c_{W}c_{12}}{2g}$
$d_{3}$ $\frac{g_{3}c_{W}c_{23}}{6g}$ $\frac{g_{3}c_{W}c_{23}}{2g}$ 0 0